### v1

# Pentagram maps over rings, Grassmannians, and skewers

Abstract
The pentagram map is a discrete dynamical system on planar polygons. By definition, the image of a polygon $P$ under the pentagram map is the polygon $P^{\prime}$ whose vertices are the intersection points of consecutive shortest diagonals of $P$. The pentagram map was introduced by R. Schwartz in 1992, and is now one of the most renowned discrete integrable systems.
Several authors proposed generalizations of the pentagram map to other geometries, in particular to Grassmannians, where the role of points and lines is played by higher-dimensional subspaces, as well to skewer geometry, where both points and lines are affine lines in the three-dimensional Euclidean space.
In the present paper, we develop a common framework for these kinds of generalizations. Specifically, we show that those maps can be viewed as pentagram maps in the projective plane over an appropriate ring. In general, those rings need not be division rings or commutative. We show that the Grassmannian pentagram map corresponds to the ring of matrices, while the skewer map is the pentagram map over the ring of dual numbers. Furthermore, we prove that the pentagram map remains integrable for any stably finite ground ring $R$.
1 Introduction
The pentagram map is a discrete dynamical system on the space of planar polygons introduced in [16]. The definition of the pentagram map is illustrated in Figure 1: the image of the polygon $P$ under the pentagram map is the polygon $P^{\prime}$ whose vertices are the intersection points of consecutive shortest diagonals of $P$, i.e. diagonals connecting second-nearest vertices. In [14, 17], it was shown that the pentagram map is a completely integrable system. As of now, it is one of the most famous discrete integrable systems which features numerous connections to various areas of mathematics. In particular, the pentagram map has an interpretation in terms of cluster algebras [7], networks of surfaces [6], the dimer model [1], Poisson-Lie groups [5], and difference operators [8]. Furthermore, the pentagram map can be viewed as a space-time discretization of the Boussinesq equation [14], a shallow water approximation.
$P$$P^{\prime}$
Figure 1: The pentagram map.
The pentagram map has been generalized in several different directions. One can intersect diagonals joining vertices that are further apart, consider polygons in higher-dimensional spaces, etc. [10, 11, 6]. Here, we are interested in different kinds of generalizations, which are based on replacing conventional notions of points and lines by more general objects with similar properties. One example of such a map is the Grassmann pentagram map, defined in [4] and further studied in [13, 2]. In this setting, the role of points is played by elements of $\mathrm{Gr}(k,3k)$, i.e. $k$-dimensional subspaces in a $3k$-dimensional vector space, while lines are elements of $\mathrm{Gr}(2k,3k)$. Any two generic points $V_{1},V_{2}\in\mathrm{Gr}(k,3k)$ belong to a single line $V_{1}\oplus V_{2}\in\mathrm{Gr}(2k,3k)$, and any two generic lines $W_{1},W_{2}\in\mathrm{Gr}(2k,3k)$ meet at a point $W_{1}\cap W_{2}\in\mathrm{Gr}(k,3k)$. The Grassmann pentagram map takes a (cyclically labeled) sequence of points $V_{i}\in\mathrm{Gr}(k,3k)$, which we think of as a polygon is the Grassmannian $\mathrm{Gr}(k,3k)$, to a new sequence given by $(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3})$. In other words, just like the classical pentagram map, the Grassmann map is defined by intersecting diagonals connecting second-nearest vertices. Furthermore, the classical pentagram map is nothing else but the $k=1$ version of the Grassmann map.
Another generalization of the pentagram map, similar in spirit, is the skewer pentagram map [20]. In this case, both points and lines are affine lines in the Euclidean space $\mathbb{R}^{3}$, with a point being incident to a line if and only if the corresponding affine lines meet a right angle. Thus, both the operation of constructing a line through two points and finding the intersection of two lines reduce to taking the common perpendicular, or skewer. Therefore, the corresponding pentagram map takes a cyclically labeled sequence $\ell_{i}$ of affine lines in $\mathbb{R}^{3}$ to the sequence $S(S(\ell_{i},\ell_{i+2}),S(\ell_{i+1},\ell_{i+3}))$, where $S(\ell,\ell^{\prime})$ stands for the skewer of lines $\ell,\ell^{\prime}$.
In the present paper, we develop a common framework for these types of pentagram-like maps. Specifically, we show that such maps can be viewed as pentagram maps in the projective plane over an appropriate ring. In general, those rings need not be division rings or commutative. Specifically, we show that the Grassmann pentagram map corresponds to the ring of matrices, while the skewer map is the pentagram map over the ring $\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ of dual numbers.
Furthermore, we prove that the pentagram map is integrable for any stably finite ground ring $R$, thus generalizing the results of [4, 13] on integrability of the Grassmann map and answering the question of [20] regarding integrability of the skewer map. Integrability is understood in this paper as the existence of a Lax form. A discrete dynamical system $\phi$ on a set $X$ (i.e., a map $\phi\colon X\to X$) is said to admit a Lax form if to every $x\in X$ one can assign an operator in such a way that operators assigned to $x$ and $\phi(x)$ are similar. In the setting of pentagram maps over a ring $R$, these operators are pseudo-difference operators with coefficients in $R$.
The existence of a Lax form is one of the characteristic features of integrable dynamical systems. In particular, for the classical pentagram map (i.e., for the pentagram map over a field), the Lax form has been a central tool for establishing the map’s dynamical properties, such as quasi-periodicity, cf. [17, 24]. The study of dynamical aspects of integrability for pentagram maps over rings is beyond the scope of the present paper. However, we do expect that quasi-periodicity-type results can be extended from fields to a fairly large class of rings, in particular to finite-dimensional algebras.
The structure of the paper is as follows.
In Section 2, we recall the definition and standard properties of stably finite rings. Details can be found, e.g., in [12, Chapter 1, §1B].
Section 3 serves as an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject (see [23]), the setting of stably finite rings appears to be new. In Section 3.1, we define projective spaces as projectivizations of arbitrary modules. The case of free modules is treated in detail in Section 3.3. We call the projectivization of a free rank $d+1$ module the projective space of dimension $d$. Sections 3.2 and 3.4 are devoted to projective spaces over the full matrix ring, which is the first of our two main examples. The second main example, the ring of dual numbers, is discussed in Section 3.5.
In Section 4, we define the notion of a polygon in the projective plane over a ring, as well as show that the pentagram map on polygons is well-defined. Furthermore, we show that, for matrix rings and dual numbers, the general notion of a pentagram map over a ring specializes to Grassmann and skewer pentagram maps respectively.
In Section 5, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators, we use that language to give an algebraic description of the moduli space of polygons in the projective plane over a ring. The main result of the section, Theorem 5.7, is a ring counterpart of [8, Proposition 3.3] which gives a description of real polygons in terms of difference operators. We note that the proof in the ring case is substantially more technical due to the lack of commutativity and essentially relies on the stable finiteness condition.
In Section 6, we establish integrability of the pentagram over an arbitrary stably finite ring. As a first step, in Section 6.1, we reformulate the pentagram map in the language of difference operators. The main result of the section is Theorem 6.5, which generalizes [8, Theorem 1.1] and essentially says that, in terms of difference operators, the pentagram map is described by a linear equation. This is used in Section 6.3 to give a Lax formulation of the pentagram map. It is given in terms of pseudo-difference operators, which are introduced in Section 6.2.
Finally, in Section 7, we use the Lax form to give invariants (i.e., conserved quantities, or first integrals) of the ring pentagram map on closed polygons. In the case of a field, these invariants are known to form a maximal Poisson-commuting family [14]. We expect that similar results can be obtained for much bigger classes of rings.
Acknowledgments. This work was supported by NSF grant DMS-2008021.
2 Stably finite rings
In this section, we recall the definition and basic properties of stably finite rings – the main class of rings relevant to the present paper; for details see [12, Chapter 1, §1B]. In what follows, all rings are assumed to be unital and non-trivial (i.e. have a multiplicative identity element $1\neq 0$), but may be non-commutative and/or have zero divisors. All modules are supposed to be unitary.
Definition 2.1
Let $R$ be a ring. An $R$-module $M$ is Hopfian if any surjective endomorphism $\phi\in\mathrm{End}_{R}(M)$ is an isomorphism.
For projective (in particular, free) modules, this condition can also be reformulated as follows.
Proposition 2.2
Suppose $R$ is a ring, and $M$ is a projective $R$-module. Then the following conditions are equivalent.
  1. 1.
    $M$ is Hopfian;
  2. 2.
    the ring $\mathrm{End}_{R}(M)$ is Dedekind finite (recall that a ring is Dedekind finite if $ab=1$ implies $ba=1$).
  3. 3.
    if $M\simeq M\oplus N$ for some other $R$-module $N$, then $N=0$.
Proof.
$1\Rightarrow 2$. Suppose $M$ is Hopfian, and let $\phi,\psi\in\mathrm{End}_{R}(M)$ be such that $\phi\psi=\mathrm{id}$. Then $\phi$ is surjective and hence an isomorphism. So, $\psi\phi=\mathrm{id}$ as well, as needed.
$2\Rightarrow 3$. Suppose the ring $\mathrm{End}_{R}(M)$ is Dedekind finite, and let $\phi\colon M\to M\oplus N$ be an isomorphism. Let also $\pi\colon M\oplus N\to M$ be the projection onto the first summand. Then, since $\pi$ is surjective, and $i\colon M\to M\oplus N$ be the inclusion map. Then $(\pi\phi)(\phi^{-1}i)=\pi i=\mathrm{id}$, so, by Dedekind finiteness, $\pi\phi$ is an isomorphism, and $N=\mathrm{Ker}\,\pi=0$.
$3\Rightarrow 1$. Suppose that $M\simeq M\oplus N$ implies $N=0$. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. Then, since $M$ is projective, the exact sequence $0\to\mathrm{Ker}\,\phi\to M\to M\to 0$ splits, and $M\simeq M\oplus\mathrm{Ker}\,\phi$. So, $\mathrm{Ker}\,\phi=0$, proving that $M$ is Hopfian. ∎
Corollary 2.3
Suppose $R$ is a ring, and $M$ is a free $R$-module of finite rank. Then $M$ is Hopfian if and only if its dual module $M^{*}=\mathrm{Hom}_{R}(M,R)$ is Hopfian.
Proof.
For a free module of finite rank, the map sending an endomorphism to its adjoint is an anti-isomorphism between the rings $\mathrm{End}_{R}(M)$ and $\mathrm{End}_{R}(M^{*})$. Since the Dedekind finiteness condition is left-right symmetric, it is preserved by anti-isomorphisms. ∎
Furthermore, if $M$ is free of finite rank (i.e., admits a finite basis), then the Hopfian condition can be restated as follows.
Proposition 2.4
Suppose an $R$-module $M$ has a basis of size $n$. Then $M$ is Hopfian if and only if any generating set of size $n$ is a basis of $M$.
Proof.
Let $v_{1},\dots,v_{n}$ be a basis of $M$. Suppose $M$ is Hopfian, and let $w_{1},\dots,w_{n}$ be a generating set. Then the endomorphism $M\to M$ taking $v_{i}$ to $w_{i}$ is surjective and hence an isomorphism. Therefore, $w_{1},\dots,w_{n}$ is a basis.
Conversely, assume that any generating set of size $n$ is a basis of $M$, and let $\phi\in\mathrm{End}_{R}(M)$ be surjective. Then $\phi$ takes $v_{1},\dots,v_{n}$ to a generating set, which, by our assumption, must be a basis again. So, $\phi$ is an isomorphism. ∎
Definition 2.5
A ring $R$ is called stably finite (or weakly finite) if any free finite rank $R$-module is Hopfian.
Since the dual of a left module is a right module and vice versa, by Corollary 2.3 we see that all free finite rank left $R$-modules are Hopfian if and only if all free finite rank right $R$-modules are Hopfian. So, Definition 2.5 might be phrased in terms of left modules or right modules.
One of the most important properties of stably finite rings is that they have an invariant basis number property:
Proposition 2.6
Assume that $R$ is stably finite, $M$ is an $R$-module, and $v_{1},\dots,v_{n}$, $w_{1},\dots,w_{m}$ are two bases in $M$. Then $m=n$.
Proof.
This is equivalent to saying that $R^{n}\simeq R^{m}$ (as left or right modules) implies $m=n$. Assume that $\phi\colon R^{n}\to R^{m}$ is an isomorphism. Without loss of generality, let $n\leq m$. Then the projection map $\pi\colon R^{m}\to R^{n}$, $\pi(r_{1},\dots,r_{m}):=(r_{1},\dots,r_{n})$ is surjective, and so is the composition $\pi\phi\colon R^{n}\to R^{n}$. Therefore, $\pi\phi$ is an isomorphism, and $\pi$ is injective. So, $m=n$. ∎
So, for every finite rank free module over a stably finite ring, the notion of rank is well-defined.
The following gives another equivalent characterization of stably finite rings.
Proposition 2.7
A ring $R$ is stably finite if and only if all matrix rings $\text{M}_{n}(R)$ are Dedekind-finite. In particular, all stably finite rings are Dedekind-finite.
Proof.
The first claim follows from Proposition 2.2, combined with the isomorphism $\mathrm{End}_{R}(M)\simeq\text{M}_{n}(R)$ for a free rank $n$ right module $M$. The second claim follows from the identification $\text{M}_{1}(R)\simeq R$. ∎
The following says that all rings that are “not too large” are stably finite.
Proposition 2.8
All Noetherian rings are stably finite.
Proof.
Let $M$ be a finite rank free module over a Noetherian ring $R$. Then $M$ is a Noetherian module. Let $\phi\in\mathrm{End}_{R}(M)$ be surjective. By the Noetherian property, the chain of submodules $\mathrm{Ker}\,\phi\subset\mathrm{Ker}\,\phi^{2}\subset\dots$ must stabilize: $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$ for some $n$. Suppose $v\in\mathrm{Ker}\,\phi$. By surjectivity, there is $w\in M$ such that $v=\phi^{n}(w)$. Then $\phi^{n+1}(w)=\phi(v)=0$, so $w\in\mathrm{Ker}\,\phi^{n+1}$. But $\mathrm{Ker}\,\phi^{n}=\mathrm{Ker}\,\phi^{n+1}$, so $v=\phi^{n}(w)=0$. Thus, $\phi$ is an isomorphism. ∎
As a corollary, all Artinian rings are stably finite. This can also be proved directly, as follows. Assume $R$ is Artinian, and let $M$ be a free finite rank module over $R$. Then $M$ is also Artinian and hence can not have proper submodules isomorphic to $M$ (otherwise, repetition gives an infinite descending chain of submodules). So, $M\simeq M\oplus N$ implies $N=0$, meaning that $R$ is stably finite.
3 Projective spaces over rings
This section is an introduction into projective geometry over rings. We note that while ring geometry is a well-studied subject (see [23]), the setting of stably finite rings appears to be new.
3.1 The projectivization of a module
There exist several, generally inequivalent, ways to define projective spaces over rings. Here we adapt the definition given in [22]. Although we will mainly be interested in projective spaces obtained from free modules, we first define the projectivization of an arbitrary module.
Definition 3.1
Let $M$ be a (left or right) $R$-module. A subspace $N\subset M$ is a free submodule which admits a direct complement in $M$ (i.e. there is a submodule $K\subset M$, not necessarily free, such that $M=N\oplus K$).
Proposition 3.2
Suppose $M$ is an $R$-module, and $N\subset M$ is a subspace. Then $N$ is also a subspace in any submodule of $M$ containing $N$.
Proof.
If $M=N\oplus K$, and $L\supset N$ is another submodule of $M$, then $L=N\oplus(K\cap L)$. ∎
Definition 3.3
The projectivization $\mathbb{P}(M)$ is the set of rank one subspaces of $M$.
One can projectivize both left modules and right modules.
Definition 3.4
Given $R$-modules $M,N$ of the same kind (both left or both right), a projective transformation $\mathbb{P}(M)\to\mathbb{P}(N)$ is a map induced by a module isomorphism $M\to N$.
For a module $M$, its projectivization $\mathbb{P}(M)$ may also be defined as the set of unimodular elements modulo scalars. For any module $M$, let $M^{*}=\mathrm{Hom}_{R}(M,R)$ be its dual module.
Definition 3.5
Let $M$ be an $R$-module. An element $v\in M$ is called unimodular if there exists $\xi\in M^{*}$ such that $\xi(v)=1$ (equivalently, if there exists $\xi\in M^{*}$ such that $\xi(v)$ is a unit).
Proposition 3.6
Let $M$ be an $R$-module. Then rank one subspaces of $M$ are precisely cyclic submodules of $M$ generated by unimodular elements.
Proof.
We give a proof for the case of left modules, the case of right modules is analogous. Suppose $P\subset M$ is a rank one subspace. Then, by definition, $P$ is freely generated over $R$ by some $v\in M$, and is a direct summand in $M$, i.e. $M=P\oplus N$. Therefore, for each $w\in M$ there exists unique $r\in R$ and $u\in N$ such that $w=rv+u$. Define a homomorphism $\xi\colon M\to R$ by $\xi(rv+u):=r$. Then $\xi(v)=1$, ensuring $v$ is unimodular.
Now, suppose that $P=Rv$ for some unimodular $v\in M$. Since $v$ is unimodular, there exists some $\xi\in M^{*}$ such that $\xi(v)=1$. Then $\xi(rv)=r$, so $rv\neq 0$ as long as $r\neq 0$, meaning that $P$ is free. Furthermore, we have $M=P\oplus\mathrm{Ker}\,\xi$, so $P$ is indeed a rank one subspace. ∎
Proposition 3.7
Let $M$ be an $R$-module. Two unimodular elements $v,w\in M$ generate the same subspace if and only if they are unit multiples of each other.
Proof.
We give a proof for the case of left modules, the case of right modules is analogous. Clearly, if $r$ is a unit, then $Rv=Rrv$. Conversely, assume that $Rv=Rw$. Then $w\in Rv$, so $w=rv$ for some $r\in R$. Likewise, $v=r^{\prime}w$ for some $r^{\prime}\in R$. So, $w=rv=rr^{\prime}w$. Since $w$ is unimodular, this implies $rr^{\prime}=1$. Similarly, $r^{\prime}r=1$. So, $r$ is a unit. ∎
Corollary 3.8
Let $M$ be an $R$-module. There is a one-to-one correspondence between points of $\mathbb{P}(M)$ and unimodular elements in $M$ modulo units.
3.2 Projective spaces over the full matrix ring
Here we consider our first main example – the projective space over the full matrix ring, cf. [23, Section 7] and [21].
Let $\mathbb{F}$ be a field, and $R=\text{M}_{k}(\mathbb{F})$ be the ring of $k\times k$ matrices over $\mathbb{F}$. By definition, projective spaces over $R$ are projectivizations of $R$-modules. Since $R$ is isomorphic to its opposite ring via transposition, there is no distinction between left and right modules. So, we will restrict our attention to left modules. Any left module $M$ over $\text{M}_{k}(\mathbb{F})$ is isomorphic to the one of the form $V^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}V$, where $V$ is a (not necessarily finite-dimensional) vector space over $\mathbb{F}$, and $\text{M}_{k}(\mathbb{F})$ acts on the first tensor factor. The correspondence $V\mapsto V^{k}$ provides additive equivalence between the categories of $\mathbb{F}$-modules (vector spaces) and $\text{M}_{k}(\mathbb{F})$-modules. This is also known as the Morita equivalence between the corresponding rings $\mathbb{F}$ and $\text{M}_{k}(\mathbb{F})$. Thus, for the purpose of describing projective spaces over $\text{M}_{k}(\mathbb{F})$, it suffices to consider left modules of the form $M=V^{k}$, where $V$ is a vector space over $\mathbb{F}$.
Proposition 3.9
There is a natural identification $\mathbb{P}(V^{k})\simeq\mathrm{Gr}(k,V)$.
Proof.
Submodules of the module $V^{k}$ are of the form $W^{k}=\mathbb{F}^{k}\otimes_{\mathbb{F}}W$ where $W\subset V$ is an arbitrary subspace of $V$. The correspondence $W\mapsto W^{k}$ gives an isomorphism between the lattice of subspaces of $V$ and the lattice of submodules of $V^{k}$. In particular, any submodule of $V^{k}$ admits a direct complement, which means that subspaces of $V^{k}$ are precisely its free submodules. Further, $W^{k}$ is a free rank one $\text{M}_{k}(\mathbb{F})$-module precisely when $\dim_{\mathbb{F}}W=k$, so rank one free submodules of $V^{k}$ are precisely of the form $W^{k}$ for $W\in\mathrm{Gr}(k,V)$. ∎
Remark 3.10
By Corollary 3.8, the space $\mathbb{P}(M)$ can alternatively be described as the set of unimodular elements in $M$ modulo units. For a vector space $V$ of finite dimension $n$, the module $V^{k}$ is isomorphic to the module $\text{M}_{k,n}(\mathbb{F})$ of $k\times n$ matrices. Such a matrix is unimodular in the sense of Definition 3.5 if and only if it has full rank. Thus
$\mathbb{P}(V^{k})=\{\mbox{full rank}\,k\times n\mbox{ matrices}\}/\{\mbox{invertible}\,k\times k\mbox{ matrices}\},$
which is the standard description of $\mathrm{Gr}(k,n)$.
3.3 The $d$-dimensional projective space
Definition 3.11
Suppose that $M$ is (left or right) free $R$-module of rank $d+1$. Then we call $\mathbb{P}(M)$ a (respectively, left or right) projective space of dimension $d$.
We have two kinds of projective spaces: left projective spaces (projectivizations of left modules) and right projective spaces (projectivizations of right modules). In the rest of this section, we restrict our attention to left projective spaces. The case of right projective spaces is considered analogously. Clearly, all (left) projective spaces of the same (finite) dimension are projectively equivalent (i.e., are related to each other by a projective transformation). We will frequently not distinguish between projectively equivalent objects and talk about the $d$-dimensional projective space over $R$, denoted $\mathbb{P}^{d}(R)$. Without loss of generality, one can think of $\mathbb{P}^{d}(R)$ as the projectivization of $R^{d+1}$.
In addition to points in projective spaces $\mathbb{P}(M)$, we will be interested in higher-dimensional subspaces, corresponding to subspaces of $M$ of rank $k\geq 2$. For stably finite rings, one has the following:
Proposition 3.12
If $M$ is a free finite rank module over a stably finite ring $R$, and $N\subsetneq M$ is a proper subspace, then $\text{rank}(N)<\text{rank}(M)$.
Proof.
Assume $N\subsetneq M$ is a subpace. If $\text{rank}(N)\geq\text{rank}(M)$ (including if $N$ has an infinite rank), then $N$ has a submodule isomorphic to $M$ which is a direct summand in $N$ and hence in $M$. But this is impossible since $M$ is Hopfian (see Proposition 2.2). ∎
Corollary 3.13
Let $R$ be a stably finite ring, and $M$ be an $R$-module. If $L\subsetneq N$ are subspaces of $M$, then $\text{rank}(L)<\text{rank}(N)$.
Proof.
By Proposition 3.2, we have that $L$ is a subspace of $N$, so the result follows from Proposition 3.12. ∎
In particular, subspaces of the projective plane $\mathbb{P}^{2}(R)$ over a stably finite ring are its points (i.e., rank one subspaces in a free rank three $R$-module $M$) and lines (i.e., rank two subspaces $L\subset M$). No point is contained in another point, and no line is contained in another line. So, the poset of subspaces of $\mathbb{P}^{2}(R)$ is determined by the incidence of points and lines. Such a structure is known as an incidence structure.
Definition 3.14
An incidence structure is a set of points, denoted by $\mathcal{P}$, a set of lines, denoted $\mathcal{L}$, and an incidence $\mathcal{I}\subset\mathcal{P}\times\mathcal{L}$. We say $P\in\mathcal{P},L\in\mathcal{L}$ are incident and write $P\in L$ if $(P,L)\in\mathcal{I}$. Two incidence structures $(\mathcal{P},\mathcal{L},\mathcal{I})$ and $(\mathcal{P}^{\prime},\mathcal{L}^{\prime},\mathcal{I}^{\prime})$ are isomorphic if there are bijections $\mathcal{P}\to\mathcal{P}^{\prime}$, $\mathcal{L}\to\mathcal{L}^{\prime}$ such that the image of $\mathcal{I}$ under the induced map $\mathcal{P}\times\mathcal{L}\to\mathcal{P}^{\prime}\times\mathcal{L}^{\prime}$ is precisely $\mathcal{I}^{\prime}$.
If $\mathcal{P}$ is a point set of some (fixed) incidence structure, we will say that $\mathcal{P}$ is an incidence space. We will say that a map $\mathcal{P}\to\mathcal{P}^{\prime}$ is an isomorphism of incidence spaces if it extends to an isomorphism of incidence structures.
We may now formalize how the projective plane $\mathbb{P}^{2}(R)=\mathbb{P}(M)$, where $M$ is a rank three free $R$-module, is an incidence space: the set of lines is
$\mathcal{L}(\mathbb{P}(M)):=\{\mbox{rank two subspaces of $M$}\},$
and the incidence $\mathcal{I}\subset\mathbb{P}(M)\times\mathcal{L}(\mathbb{P}(M))$ is given by containment ($(P,L)\in\mathcal{I}\iff P\subset L$).
Clearly, any projective transformation $\mathbb{P}(M)\to\mathbb{P}(N)$, where $M,N$ are rank three free $R$-modules, is also an isomorphism of incidence spaces. Thus, all (left) projective planes over $R$ are isomorphic as incidence spaces, and one can talk about the natural incidence structure on $\mathbb{P}^{2}(R)$. In a similar way, one can define higher rank incidence structures on projective spaces $\mathbb{P}^{d}(R)$, $d>2$.
3.4 The projective plane over the full matrix ring
Back to the example $R=\text{M}_{k}(\mathbb{F})$, notice that $V^{k}$ is a free $\text{M}_{k}(\mathbb{F})$-module of rank $d+1$ if and only if $\dim V=k(d+1)$. Thus, by Proposition 3.9, we have
$\mathbb{P}^{d}(\text{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,(d+1)k).$
In particular,
$\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathrm{Gr}(k,3k).$
Furthermore, the space of lines in $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows.
Proposition 3.15
There is a natural identification $\mathcal{L}(\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F})))\simeq\mathrm{Gr}(2k,3k)$. The incidence relation between points $P\in\mathrm{Gr}(k,3k)$ and lines $L\in\mathrm{Gr}(2k,3k)$ is given by containment.
Proof.
We have $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. Recall that the correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with lattice of submodules of $V^{k}$. Furthermore, $W^{k}$ is free of rank two if and only if $\dim W=2k$. Thus, lines in $\mathbb{P}(V^{k})$ correspond to $2k$-dimensional subspaces of $V$, and incidence between points and lines is given by containment. ∎
In words, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ can be described as follows: points are $k$-dimensional subspaces of $\mathbb{F}^{3k}$, lines are $2k$-dimensional subspaces of $\mathbb{F}^{3k}$, while incidence of points and lines corresponds to containment. Note that two generic points $P_{1},P_{2}\in\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ lie on a unique line $P_{1}\oplus P_{2}$. Likewise, two generic lines $L_{1},L_{2}$ meet at a unique point $L_{1}\cap L_{2}$. However, if $\dim P_{1}\oplus P_{2}<2k$, then there exist infinitely many lines containing points $P_{1},P_{2}$ (and, if $\dim L_{1}\cap L_{2}>k$, there exist infinitely many points lying on $L_{1},L_{2}$). Thus, $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is not a projective plane in the sense of incidence geometry. Incidence spaces with properties similar to those of $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ are known as Barbilian planes, see [3, 22]. It is shown in [22] that such planes are precisely the projective planes over rings of stable rank two, of which the matrix ring is an example.
3.5 The projective plane over the dual numbers
Let $\mathbb{D}:=\mathbb{R}[\varepsilon]/(\varepsilon^{2})$ be the ring of dual numbers. In this section, we show that the geometry of $\mathbb{P}^{2}(\mathbb{D})$ is the skewer geometry of [19], where both points and lines are affine lines in $\mathbb{R}^{3}$, and a point is incident to a line if and only if the corresponding affine lines meet at a right angle.
Let $V$ be a three-dimensional vector space over $\mathbb{R}$. Then $\mathbb{D}\otimes_{\mathbb{R}}V$ is a free rank three $\mathbb{D}$-module. Our aim is to give a geometric description of the incidence space $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$, to which end we fix a positive-definite inner product on $V$. Of course, the incidence structure on $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ does not depend on that inner product. However, our geometric model does.
Let $\mathcal{A}(V)$ be the space of affine straight lines in $V$. We define an incidence structure on $\mathcal{A}(V)$ as follows: both points and lines are $\mathcal{P}=\mathcal{L}=\mathcal{A}(V)$ and a point is incident to a line if and only if the affine lines intersect and are perpendicular to each other. In this geometry, both the operation of joining two points with a line and intersecting two lines boil down to taking the common perpendicular, or skewer, of the corresponding affine lines.
We will show that there is natural isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$. The corresponding mapping $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ between the point spaces is a variant of the Study mapping, see [15, Section 2.3] and [18]. For an affine line $\{p+tv\mid t\in\mathbb{R}\}$, where $p,v\in V$ and $v\neq 0$, let
$\zeta(v,p):=v+\varepsilon v\times p,$ (3.1)
where the notation $\times$ stands the cross product in $V$, and define
$\phi_{P}(p+tv):=\mathbb{D}\zeta(v,p).$
Proposition 3.16
The mapping $\phi_{P}$ is well-defined and bijective.
Proof.
First, note that a dual vector $v+\varepsilon w\in\mathbb{D}\otimes_{\mathbb{R}}V$ (where $v,w\in V$) is unimodular if and only if $v\neq 0$. Indeed, if $v\neq 0$, then, for the $\mathbb{D}$-linear extension of any $\xi\in V^{*}$ such that $\xi(v)\neq 0$, we have that $\xi(v+\varepsilon w)$ is a unit. Conversely, if $v=0$, then for any $\xi\in(\mathbb{D}\otimes_{\mathbb{R}}V)^{*}$, we have that $\xi(v+\varepsilon w)$ is divisible by $\varepsilon$ and thus is not a unit.
In particular, for any $v\neq 0$ we have that $\zeta(v,p)$ is unimodular and therefore $\mathbb{D}\zeta(v,p)$ is a subspace. Moreover, this subspace is independent on the choice of the direction vector $v$ and base point $p$ of the line $v+tp$, since for any $\lambda\in\mathbb{R}$ we have $\zeta(\lambda v,p)=\lambda\zeta(v,p)$ and $\zeta(v,p+\lambda v)=\zeta(v,p)$. So, the mapping $\phi_{P}$ is well-defined.
Verification of bijectivity is straightforward. ∎
Proposition 3.17
Let $M$ be a finite rank free $\mathbb{D}$-module, and suppose that $M=N\oplus K$. Then $N$ and $K$ are free modules.
Proof.
A $\mathbb{D}$-module is an $\mathbb{R}[\varepsilon]$-module annihilated by $\varepsilon^{2}$. Therefore, any finitely generated $\mathbb{D}$-module $M$ is isomorphic to a unique module of the form $M^{p,q}=\mathbb{D}^{\oplus p}\oplus(\mathbb{D}/(\varepsilon))^{\oplus q}$ (this is just the Jordan decomposition over the reals for the multiplication by $\varepsilon$ in $M$). Such a module is free if and only if $q=0$. So, the direct sum of two modules, at least one of which is not free, cannot be free. ∎
Proposition 3.18
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace of rank $k$. Then $N^{\bot}:=\{v\in M\mid\langle v,w\rangle=0\mbox{ for all }w\in N\}$ is also a subspace of rank $n-k$.
Proof.
Consider the restriction mapping $\rho\colon M^{*}\to N^{*}$. Since $N$ is free of rank $k$, so is $N^{*}$, and $M^{*}=\mathrm{Ker}\,\rho\oplus K$ for some submodule $K\subset M^{*}$. Thus, by Proposition 3.17, the submodule $\mathrm{Ker}\,\rho\subset M$ is free. Furthermore, $\rho_{K}\colon K\to N^{*}$ is an isomorphism, so $K$ has rank $k$ and $\mathrm{Ker}\,\rho$ has rank $n-k$. Now, consider the mapping $\phi\colon M\to M^{*}$ given by the inner product. Since the inner product is non-degenerate, the mapping $\phi$ is injective. Therefore, since $M$ is a finite-dimensional real algebra, $\phi$ is, in fact, an isomorphism. Together, $N^{\bot}=\phi^{-1}(\mathrm{Ker}\,\rho)$ is a free submodule of $M$ which has rank $n-k$ and admits a direct complement, $\phi^{-1}(K)$, as needed. ∎
Corollary 3.19
Suppose $M$ is a free $\mathbb{D}$-module of finite rank, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Let $N\subset M$ be a subspace. Then $(N^{\bot})^{\bot}=N$.
Proof.
We have $(N^{\bot})^{\bot}\supset N$. At the same time, by Proposition 3.18, the subspaces $(N^{\bot})^{\bot}$ and $N$ are of the same (finite) rank. Therefore, by Corollary 3.13, we have $(N^{\bot})^{\bot}=N$. (Note that $\mathbb{D}$ is a finite-dimensional real algebra, hence Noetherian, hence stably finite.) ∎
Corollary 3.20
Suppose $M$ is a free $\mathbb{D}$-module of rank $n$, equipped with a non-degenerate $\mathbb{D}$-linear inner product. Then the correspondence $N\mapsto N^{\bot}$ is a bijection:
{subspaces of $\{\text{subspaces of }M\mbox{of rank }k\}\leftrightarrow\{\text{subspaces of }M\mbox{of rank }n-k\}.$
We now define a map $\phi_{L}\colon\mathcal{A}(V)\to\mathcal{L}(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ by
$\phi_{L}(p+tv):=\phi_{P}(p+tv)^{\bot},$
where the inner product on $\mathbb{D}\otimes_{\mathbb{R}}V$ is defined as the unique $\mathbb{D}$-linear extension of the inner product on $V$. It follows from Proposition 3.16 and Corollary 3.20 that the map $\phi_{L}$ is well-defined and bijective.
Proposition 3.21
The mappings $\phi_{P}\colon\mathcal{A}(V)\to\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)$ and $\phi_{L}\colon\mathcal{A}(V)\to\mathcal{L}(\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V))$ provide an isomorphism of incidence spaces $\mathbb{P}(\mathbb{D}\otimes_{\mathbb{R}}V)\simeq\mathcal{A}(V)$.
Proof.
We already know that the mappings $\phi_{P}$, $\phi_{L}$ are bijections, so it remains to show that they preserve incidence. In other words, we need to show that the lines $p+tv$ and $q+tw$ meet at a a right angle if and only if $\phi_{P}(p+tv)\subset\phi_{L}(q+tw)$. We have
$\phi_{P}(p+tv)\subset\phi_{L}(q+tw)=\phi_{P}(q+tw)^{\bot}\quad\Leftrightarrow\quad\langle\zeta(p+tv),\zeta(q+tw)\rangle=0.$
Further,
$\langle\zeta(p+tv),\zeta(q+tw)\rangle=\langle v+\varepsilon v\times p,w+\varepsilon w\times q\rangle=\langle v,w\rangle+\varepsilon\det(v,w,q-p).$
But $\det(v,w,q-p)=0$ if either $v$ and $w$ are collinear (which can not happen if $\langle v,w\rangle=0$, since the inner product on $V$ is positive-definite), or $q-p\in\text{span}\langle v,w\rangle$. The latter is equivalent to saying that the lines $p+tv$ and $q+tw$ have a point in common. So, combining this with the condition $\langle v,w\rangle=0$, we get that $\phi_{P}(p+tv)\subset\phi_{L}(q+tw)$ if and only if the lines $p+tv$ and $q+tw$ meet at a right angle, as needed. ∎
Remark 3.22
We note that, when restricted to lines through the origin, the map $\phi_{P}$ is essentially the identity map on $\mathbb{R}\mathbb{P}^{2}$, as it sends such a line to its direction; as for $\phi_{L}$, it is the polar duality map sending a line to its orthogonal plane. In this case, Proposition 3.21 just says that points $P_{1},P_{2}\in\mathbb{R}\mathbb{P}^{2}$ are polar if and only if $P_{1}$ belongs to the polar line of $P_{2}$.
4 Polygons and pentagram maps over rings
In this section, we define the notion of a polygon in the projective plane over a ring (see Definition 4.2), as well as show that the pentagram map on polygons is well-defined (see Section 4.2). Furthermore, we show that, for matrix rings and dual numbers, the general notion of a pentagram map over a ring specializes to Grassmann and skewer pentagram maps respectively (Examples 4.8 and 4.10).
4.1 Polygons in projective spaces over rings
Over a field, the pentagram map is well-defined on generic polygons. In particular, one can always apply the pentagram to a polygon such that all consecutive triple of vertices are in general position. The general position requirement means that the points are not collinear. In the context of (stably finite) rings, the right condition that ensures that the pentagram map is defined is the spanning condition.
Definition 4.1
Let $M$ be an $R$-module. We say that points $P_{1},\dots,P_{n}\in\mathbb{P}(M)$ span $\mathbb{P}(M)$ if $P_{1}+\dots+P_{n}=M$.
By Proposition 2.4, if $R$ is stably finite, and $M$ is free of rank $n$, then $n$ points $P_{1},\dots,P_{n}$ span $\mathbb{P}(M)$ if and only if $M=P_{1}\oplus\dots\oplus P_{n}$. In other words, if $v_{i}$ spans $P_{i}$, then $P_{1},\dots,P_{n}$ span $\mathbb{P}(M)$ if and only if $v_{1},\dots,v_{n}$ is a basis of $M$.
Definition 4.2
A polygon in $\mathbb{P}^{d}(R)$ is a bi-infinite sequence of points $P_{i}\in\mathbb{P}^{d}(R)$ such that, for any $i\in\mathbb{Z}$, the points $P_{i},P_{i+1},\dots,P_{i+d}$ span $\mathbb{P}^{d}(R)$.
For stably finite $R$, this definition can be restated as follows: if $\mathbb{P}^{d}(R)=\mathbb{P}(M)$, and $v_{i}$ spans $P_{i}$, then $v_{i},v_{i+1},\dots,v_{i+d}$ is a basis of $M$ for any $i\in\mathbb{Z}$.
We notice that our definition of polygons allows either infinite polygons or closed ones, i.e., polygons such that $P_{i}=P_{i+n}$ for some fixed $n\in\mathbb{Z}_{+}$ and all $i\in\mathbb{Z}$.
Definition 4.3
We say two polygons $(P_{i})$ and $(P_{i}^{\prime})$ are projectively equivalent if there exists a fixed projective transformation that takes each $P_{i}$ to $P_{i}^{\prime}$.
For most of the paper, we are interested in polygons considered up to projective equivalence.
Remark 4.4
We note that the notion of being spanning, and hence the notion of a polygon, is not determined by the incidence structure. For example, it is easy to see that the inclusion map $\mathbb{Z}^{3}\to\mathbb{Q}^{3}$ gives an isomorphism of incidence spaces $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ (the same is true for any PID and its field of fractions). However, while any three non-collinear points $P_{1},P_{2},P_{3}$ span $\mathbb{P}^{2}(\mathbb{Q})$, for $\mathbb{P}^{2}(\mathbb{Z})$ that is only so if the determinant of the matrix formed by the basis vectors in $P_{1},P_{2},P_{3}$ is ±$\pm 1$. Accordingly, not any polygon over $\mathbb{Q}$ is a polygon over $\mathbb{Z}$ in the sense of Definition 4.2, even though $\mathbb{P}^{2}(\mathbb{Z})\simeq\mathbb{P}^{2}(\mathbb{Q})$ as incidence spaces.
4.2 Pentagram maps over stably finite rings, Grassmannians, and skewers
Definition 4.5
Suppose $R$ is a stably finite ring. The pentagram map in $\mathbb{P}^{2}(R)$ takes a polygon $P_{i}\in\mathbb{P}^{2}(R)$ to the sequence of points $P^{\prime}_{i}\in\mathbb{P}^{2}(R)$, where $P_{i}^{\prime}$ is the intersection of the lines connecting $P_{i}$ to $P_{i+2}$ and $P_{i+1}$ to $P_{i+3}$.
The sequence $P_{i}^{\prime}$ is well-defined thanks to the following.
Lemma 4.6
Suppose $R$ is a stably finite ring, and $M$ is a free rank three $R$-module. Let $P_{1},P_{2},P_{3},P_{4}\in\mathbb{P}(M)$ be such that the triples $P_{1},P_{2},P_{3}$ and $P_{2},P_{3},P_{4}$ span $\mathbb{P}(M)$. Then
  1. 1.
    For each of the pairs of points $(P_{1},P_{2})$, $(P_{3},P_{4})$, there is a unique line containing that pair of points.
  2. 2.
    Those lines $P_{1}P_{2}$ and $P_{3}P_{4}$ meet at a single point.
Proof.
To prove the first claim, note that since $P_{1},P_{2},P_{3}$ span $\mathbb{P}(M)$, we have that $P_{1}\oplus P_{2}$ is free, and $(P_{1}\oplus P_{2})\oplus P_{3}=M$, so $P_{1}\oplus P_{2}\subset M$ is a rank two subspace (i.e., a line in $\mathbb{P}(M)$). Furthermore, any other line containing $P_{1},P_{2}$ must contain $P_{1}\oplus P_{2}$ and thus coincide with it by Corollary 3.13. So, $P_{1},P_{2}$ lie on a unique line, $P_{1}\oplus P_{2}$. Likewise, $P_{3},P_{4}$ also belong to a unique line, $P_{3}\oplus P_{4}$.
To prove the second claim, it suffices to show that $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ is a rank one subspace (i.e., a point in $\mathbb{P}(M)$). If this is done, then, by Corollary 3.13, any other point contained in the lines $P_{1}\oplus P_{2}$, $P_{3}\oplus P_{4}$ must coincide with $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$.
First, observe that since $M=P_{2}\oplus(P_{3}\oplus P_{4})$, and $P_{2}\subset P_{1}\oplus P_{2}$, we have $P_{1}\oplus P_{2}=P_{2}\oplus((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4}))$, and hence
$M=(P_{2}\oplus P_{3})\oplus((P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})).$
So, the submodule $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ admits a direct complement. Further, we also have
$M=(P_{2}\oplus P_{3})\oplus P_{1},$
so $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})\simeq P_{1}$, with the isomorphism given by the projection along $P_{2}\oplus P_{3}$. Thus, $(P_{1}\oplus P_{2})\cap(P_{3}\oplus P_{4})$ is indeed a subspace of rank one, as needed. ∎
Applying Lemma 4.6 to vertices $P_{i},P_{i+2},P_{i+1},P_{i+3}$ (in that specific order) of a polygon $P_{i}$, we get that the pentagram map on polygons is well-defined. We note that the pentagram map is not a map from the set of polygons to itself, since the image of a polygon under the pentagram map is not necessarily a polygon. Instead, it should be understood as a binary relation on the set of polygons. In the case when the base ring is a field, the pentagram map is, actually, a rational mapping.
Applying Lemma 4.6 to vertices $P_{i},P_{i+1},P_{i+2},P_{i+3}$ (in that order) of a polygon $P_{i}$, one also gets the following map:
Definition 4.7
The inverse pentagram map in $\mathbb{P}^{2}(R)$ takes a polygon $P_{i}\in\mathbb{P}^{2}(R)$ to the sequence of points $P^{\prime}_{i}\in\mathbb{P}^{2}(R)$, where $P_{i}^{\prime}$ is the intersection of the lines connecting $P_{i}$ to $P_{i+1}$ and $P_{i+2}$ to $P_{i+3}$.
This is indeed the inverse of the pentagram map in the sense that the composition of this map and the pentagram map (taken in either order) is the identity, up to shifting indices, wherever that composition is defined.
We note that both the pentagram map and its inverse commute with projective transformation and, as such, are well-defined on the set of polygons modulo projective equivalence. Also, note that the image of a closed polygon under the pentagram map is again a closed polygon.
Example 4.8
The pentagram map over the matrix ring $\text{M}_{k}(\mathbb{F})$ is precisely (the planar case of) the Grassmann pentagram map, as defined in [4]:
Proposition 4.9
A polygon in $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))$ is a sequence of $k$-dimensional subspaces $V_{i}\subset\mathbb{F}^{3k}$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=\mathbb{F}^{3k}$. The pentagram map send such a sequence to the sequence
$V_{i}^{\prime}:=(V_{i}\oplus V_{i+2})\cap(V_{i+1}\oplus V_{i+3}).$ (4.1)
Proof.
We have $\mathbb{P}^{2}(\text{M}_{k}(\mathbb{F}))=\mathbb{P}(V^{k})$, where $V$ is a vector space over $\mathbb{F}$ of dimension $3k$. The correspondence $W\mapsto W^{k}$ identifies the lattice of subspaces of $V$ with lattice of submodules of $V^{k}$. Under this correspondence, points of $\mathbb{P}(V^{k})$ are identified with $k$-dimensional subspaces of $V$. Three points span $\mathbb{P}(V^{k})$ if and only if the corresponding $k$-dimensional subspaces span $V$. Thus, a polygon in $\mathbb{P}(V^{k})$ is precisely a collection $V_{i}\in\mathrm{Gr}(k,V)$ such that $V_{i}\oplus V_{i+1}\oplus V_{i+2}=V$. Further, intersections and spans of subspaces in $V^{k}$ correspond to intersections of subspaces in $V$, giving the claimed formula for the pentagram map. ∎
Example 4.10
The pentagram map over the ring $\mathbb{D}$ of dual numbers is the skewer pentagram map, as defined in [20]:
Proposition 4.11
A polygon in $\mathbb{P}^{2}(\mathbb{D})$ is a bi-infinite sequence $\ell_{i}$ of affine lines in $\mathbb{R}^{3}$ such that, for any $i\in\mathbb{Z}$, the direction vectors of the lines $\ell_{i},\ell_{i+1},\ell_{i+2}$ are linearly independent. The pentagram map sends such a sequence to the sequence
$\ell_{i}^{\prime}:=S(S(\ell_{i},\ell_{i+2}),S(\ell_{i+1},\ell_{i+3})),$
where $S(\ell,\ell^{\prime})$ stands for the common perpendicular (skewer) of the lines $\ell,\ell^{\prime}$.
Proof.
Consider three affine lines $\{p_{1}+tv_{1}\}$, $\{p_{2}+tv_{2}\}$, $\{p_{3}+tv_{3}\}$ in $\mathbb{R}^{3}$. The associated points in $\mathbb{P}^{2}(\mathbb{D})$ are the spans of the dual vectors $\zeta(v_{i},p_{i})=v_{i}+\varepsilon v_{i}\times p_{i}$. Those points span $\mathbb{P}^{2}(\mathbb{D})$ if and only if the vectors $\zeta(v_{i},p_{i})$ span $\mathbb{D}^{3}$. Since $\mathbb{D}$ is a commutative ring, this is equivalent to saying that $\det(\zeta(v_{1},p_{1}),\zeta(v_{2},p_{2}),\zeta(v_{3},p_{3}))$ is a unit. Consider the real part function $\Re\colon\mathbb{D}\to\mathbb{R}$, $\Re(a+b\varepsilon):=a$. It is a homomorphism of rings, so
$\Re\det(\zeta(v_{1},p_{1}),\zeta(v_{2},p_{2}),\zeta(v_{3},p_{3}))=\det(\Re\zeta(v_{1},p_{1}),\Re\zeta(v_{2},p_{2}),\Re\zeta(v_{3},p_{3}))=\det(v_{1},v_{2},v_{3}).$
Therefore, three affine lines span $\mathbb{P}^{2}(\mathbb{D})$ if and only if their direction vectors span $\mathbb{R}^{3}$, giving the desired description of polygons. As for the formula for the pentagram map, it follows from the definition of the latter and Proposition 3.21. ∎
5 Polygons via difference operators
In this section, we introduce our main technical tool: difference operators with coefficients in a ring. After discussing basic properties of such operators in Section 5.1, we use that language to give an algebraic description of the moduli space of polygons in the projective plane over a ring (Theorem 5.7).
5.1 Difference operators with coefficients in a ring
For a ring $R$, denote by $R^{\mathbb{Z}}$ the ring of bi-infinite sequences of elements of $R$. We will think of $R^{\mathbb{Z}}$ as a left module over itself and, simultaneously, as a right $R$-module. Denote by $\mathrm{End}_{R}(R^{\mathbb{Z}})$ the ring of endomorphisms of $R^{\mathbb{Z}}$, viewed as a right $R$-module.
Definition 5.1
The ring of (left) difference operators over $R$ is a subring $R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with the left shift operator $T\in\mathrm{End}_{R}(R^{\mathbb{Z}})$, given by $(Ta)_{i}:=a_{i+1}$. In other words, a difference operator is an $R$-endomorphism $\mathcal{D}:R^{\mathbb{Z}}\to R^{\mathbb{Z}}$ given by
$\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i},$
where $d\geq 0$ is an integer, and $\alpha_{1},\dots,\alpha_{d}\in R^{\mathbb{Z}}$. We will denote the entries of the bi-infinite sequence $\alpha_{i}$ by $\alpha_{ij}$. For any $a\in R^{\mathbb{Z}}$, the operator $\mathcal{D}$ acts on $a$ as
$(\mathcal{D}a)_{j}=\sum_{i=0}^{d}\alpha_{ij}a_{i+j}.$
Assuming $\alpha_{d}\neq 0$, we call $d$ the degree of the difference operator $\mathcal{D}$.
More generally, one can consider the subring $R^{\mathbb{Z}}[T,T^{-1}]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with $T$ and $T^{-1}$, i.e., allow for difference operators of the form
$\mathcal{D}=\sum_{i=m}^{n}\alpha_{i}T^{i},$
where $m,n\in\mathbb{Z}$. However, we will focus primarily on operators as in Definition 5.1.
One can also define right difference operators: they are endomorphisms of $R^{\mathbb{Z}}$ viewed as a left $R$-module. A right difference operator reads
$(\mathcal{D}a)_{j}=\sum_{i=0}^{d}a_{i+j}\alpha_{ij}.$
However, for the remainder of the paper, all difference operators are assumed to be left difference operators, unless stated otherwise.
Definition 5.2
Let $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}\in\mathbb{R}^{\mathbb{Z}}[T]$ be a difference operator. If $\alpha_{0}$ and $\alpha_{d}$ are sequences of units in $R$ (i.e., are units in $R^{\mathbb{Z}}$), we call $\mathcal{D}$ a properly bounded difference operator.
Proposition 5.3
Let $\mathcal{D}\in R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ be a properly bounded difference operator of degree $d$. Then its kernel $\mathrm{Ker}\,\,\mathcal{D}\subset R^{\mathbb{Z}}$ is a free rank $d$ right $R$-module.
Proof.
Since $\mathcal{D}$ is an endomorphism of the right $R$-module $R^{\mathbb{Z}}$, its kernel is also a right $R$-module. To prove that it is free of rank $d$, notice that the equation $\mathcal{D}a=0$ may be viewed as a sequence of recurrence relations, allowing one to find all entries of $a\in R^{\mathbb{Z}}$ from $a_{1},\dots,a_{d}$. Thus, the projection mapping $\mathrm{Ker}\,\,\mathcal{D}\to R^{d}$, given by $a\mapsto(a_{1},\dots,a_{d})$, is an isomorphism of right $R$-modules, and $\mathrm{Ker}\,\,\mathcal{D}$ is free of rank $d$. ∎
If we have a difference operator $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ which is not properly bounded, there is no bound on the “size” of the kernel in terms of the degree. In particular, the kernel does not have to be free and/or of finite rank. Nevertheless, we have the following.
Proposition 5.4
Let $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$, $\mathcal{D}\neq 0$ be a difference operator of degree less than $d$. Then, for any properly bounded difference operator $\mathcal{D}^{\prime}\in\mathbb{R}^{\mathbb{Z}}[T]$ of degree $d$, there exists $a\in\mathrm{Ker}\,\mathcal{D}^{\prime}$ such that $a\notin\mathrm{Ker}\,\mathcal{D}$.
Proof.
Since $\mathcal{D}\neq 0$, there is $i\in\mathbb{Z}$ such that the condition $\mathcal{D}a=0$, where $a\in R^{\mathbb{Z}}$, imposes a non-trivial relation on $d$ consecutive entries $a_{i},\dots,a_{i+d-1}$ of $a$. On other hand, the map $\mathrm{Ker}\,\mathcal{D}^{\prime}\to R^{d}$ given by $a\mapsto(a_{i},\dots,a_{i+d-1})$ is a bijection, so we can find an element of $\mathrm{Ker}\,\mathcal{D}^{\prime}$ whose entries do not satisfy the relation forced by $\mathcal{D}a=0$. ∎
Now, let $M$ be a free finite-rank $R$-module. Then $R^{\mathbb{Z}}\otimes_{R}M\simeq M^{\mathbb{Z}}$, as left $R^{\mathbb{Z}}$-modules (where $M^{\mathbb{Z}}$ consists of bi-infinite sequences valued in $M$). Any difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ acts on $M^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}M$ by $\mathcal{D}(a\otimes v)=(\mathcal{D}a)\otimes v$, where $a\in R^{\mathbb{Z}}$, $v\in M$. Put differently, for $\mathcal{D}=\sum_{i=0}^{d}\alpha_{i}T^{i}$ and $v\in M^{\mathbb{Z}}$, we have
$(\mathcal{D}v)_{j}=\sum_{i=0}^{d}\alpha_{ij}v_{i+j}.$
Let $M$ be a free finite-rank left $R$-module with basis $e_{1},\dots,e_{m}$. For
$v=\sum_{i=1}^{m}a_{i}\otimes e_{i}\in R^{\mathbb{Z}}\otimes_{R}M=M^{\mathbb{Z}},$
define a submodule $\mathrm{Span}\,v$ of the right $R$-module $R^{\mathbb{Z}}$ as the right span of $a_{1},\dots,a_{m}$. In other words, choosing a basis in $M$ allows one to write $v\in M^{\mathbb{Z}}$ as an infinite matrix
($\ v_{i,1}&\dots&v_{i,m}\\ \vdots&&\vdots\end{array}\right)$ (5.1)
with columns indexed by basis elements and rows indexed by $\mathbb{Z}$. The rows of this matrix are just the basis expansions of the entries of the sequence $v$. Then $\mathrm{Span}\,v$ is the column space of this matrix. Note that the submodule $\mathrm{Span}\,v$ does not depend on the choice of a basis in $M$, as it can be defined as the set bi-infinite sequences of the form $(\xi(v_{i}))_{i\in\mathbb{Z}}$, where $\xi\in M^{*}$.
Proposition 5.5
Suppose $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ is a difference operator, and $v\in M^{\mathbb{Z}}$, where $M$ is a free finite-rank left $R$-module. Then $\mathcal{D}v=0$ if and only if $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$.
Proof.
Let $e_{1},\dots,e_{m}$ be a basis in $M$. Then $v=\sum_{i=1}^{m}a_{i}\otimes e_{i}$ for some $a_{i}\in R^{\mathbb{Z}}$. The submodule $\mathrm{Span}\,v$ is spanned by the $a_{i}$’s. At the same time,
$\mathcal{D}(v)=\sum_{i=1}^{m}\mathcal{D}(a_{i})\otimes e_{i},$
which is 00 if and only if $\mathcal{D}(a_{i})$ for all $i$, which is the same as to say $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$. ∎
The following proposition is a difference counterpart of the following standard result from the theory of ordinary differential equations: for functions $f_{1}(x),\dots,f_{n}(x)$ belonging to a kernel of the differential operator
$\Delta=(d/dx)^{n}+u_{n-1}(x)(d/dx)^{n-1}+\dots+u_{0},$
[…]
Proposition 5.6
Let $R$ be a stably finite ring. Suppose $\mathcal{D}\in\mathbb{R}^{\mathbb{Z}}[T]$ is a properly bounded difference operator of degree $d$, and $v\in M^{\mathbb{Z}}$, where $M$ is a free left $R$-module of rank $d$. Suppose that $\mathcal{D}v=0$. Then the following conditions are equivalent:
  1. 1.
    $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$.
  2. 2.
    $v_{k},\dots,v_{k+d-1}$ span $M$ (equivalently, form a basis in $M$) for some $k\in\mathbb{Z}$.
  3. 3.
    $v_{k},\dots,v_{k+d-1}$ span $M$ (equivalently, form a basis in $M$) for all $k\in\mathbb{Z}$.
Proof.
Let $k\in\mathbb{Z}$. It suffices to show that $v_{k},\dots,v_{k+d-1}$ span $M$ if and only if $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$. Take a basis $e_{1},\dots,e_{d}$ in $M$. Consider a $d\times d$ matrix
$V_{k}:=\left(\begin{array}[]{ccc}v_{k,1}&\dots&v_{k,d}\\ \vdots&&\vdots\\ v_{k+d-1,1}&\dots&v_{k+d-1,d}\end{array}\right)\in\text{M}_{d}(R)$
where $v_{i,j}\in R$ are coordinates of $v_{i}$ relative to the basis $e_{1},\dots,e_{d}$. The rows of the matrix $V_{k}$ are coordinate expressions for $v_{k},\dots,v_{k+d-1}\in M$, so $V_{k}$ is left invertible if and only if $v_{k},\dots,v_{k+d-1}$ span $M$. At the same time, the matrix $V_{k}$ consists of $d$ consecutive rows of the infinite matrix (5.1) whose columns span the module $\mathrm{Span}\,v$. Since taking $d$ consecutive entries gives an isomorphism $\mathrm{Ker}\,\,\mathcal{D}\simeq R^{d}$, the columns of the matrix $V_{k}$ span $R^{d}$ if and only if the columns of the matrix (5.1) span $\mathrm{Ker}\,\,\mathcal{D}$, which is equivalent to saying that $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$. To sum up, $v_{k},\dots,v_{k+d-1}$ span $M$ is and only if the matrix $V_{k}$ is left-invertible, while $\mathrm{Span}\,v=\mathrm{Ker}\,\,\mathcal{D}$ if and only if $V_{k}$ is right-invertible. Since $R$ is stably finite, those conditions are equivalent. ∎
5.2 The correspondence between difference operators and polygons
We will now establish a correspondence between polygons and difference operators.
Note that the ring $\mathbb{R}^{\mathbb{Z}}[T]$ of difference operators over $R$ contains a subring $R^{\mathbb{Z}}$ of bi-infinite sequences valued in $R$: such sequences are just degree zero difference operators. In particular, the group of units in $R^{\mathbb{Z}}$ (which consists of bi-infinite sequences of units in $R$) acts on $\mathbb{R}^{\mathbb{Z}}[T]$ in two different ways: by left multiplication, and by right multiplication.
Theorem 5.7.
Let $R$ be a stably finite ring. Then, there is a natural one-to-one correspondence between the following sets:
  1. 1.
    Projective equivalence classes of polygons in the left projective space $\mathbb{P}^{d-1}(R)$.
  2. 2.
    Properly bounded left difference operators of degree $d$ with coefficients in $R$, up to multiplication on the left and right by bi-infinite sequences of units in $R$.
By the set of projective equivalence classes of polygons in the left projective space $\mathbb{P}^{d-1}(R)$, we mean the set polygons in projectivizations of all rank $d$ free left $R$-modules, considered up to projective equivalence.
Similarly, there is a natural one-to-one correspondence between the set of projective equivalence classes of polygons in the right projective space $\mathbb{P}^{d-1}(R)$, and the set of properly bounded right difference operators of degree $d$ with coefficients in $R$, up to multiplication on the left and right by bi-infinite sequences of units in $R$. This can be proved analogously to the left case, or deduced from the left case by viewing right $R$-modules as left modules over the opposite ring.
In the case when the base ring $R$ is a field, the correspondence between polygons and difference operators is well-known. In particular, [8] gives a proof of Theorem 5.7 in the case $R=\mathbb{R}$.
In order to prove Theorem 5.7, we will first build a map from the space of all degree $d$ properly bounded left difference operators with coefficients in a stably finite ring $R$ to the space of polygons, up to projective equivalence, in the left projective space $\mathbb{P}^{d-1}(R)$. Then, we will consider difference operators up to the action bi-infinite sequences of units on left and right and show that our original map factors through this quotient. Finally, we will prove that the map on the quotient is a bijection.
Let us denote the space of all degree $d$ properly bounded difference operators over $R$ as $R^{\mathbb{Z}}[T]_{d}^{pb}$ and let us denote the space of polygons in $\mathbb{P}^{d-1}(R)$ up to projective equivalence as $\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Our first aim to build a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$. Then, by Proposition 5.3, we know that $\mathrm{Ker}\,\,\mathcal{D}\subset R^{\mathbb{Z}}$ is a rank $d$ free right $R$-module. Thus, $(\mathrm{Ker}\,{\mathcal{D}})^{*}=\mathrm{Hom}_{R}(\mathrm{Ker}\,{\mathcal{D}},R)$ is a free rank $d$ left $R$-module. For any $i\in\mathbb{Z}$, let $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,{\mathcal{D}})^{*}$ be defined by $\varepsilon_{\mathcal{D},i}(a):=a_{i}$ for any $a\in\mathrm{Ker}\,{\mathcal{D}}$. Then $\varepsilon_{\mathcal{D}}\in((\mathrm{Ker}\,{\mathcal{D}})^{*})^{\mathbb{Z}}=R^{\mathbb{Z}}\otimes_{R}(\mathrm{Ker}\,{\mathcal{D}})^{*}$.
Proposition 5.8
We have $\mathrm{Span}\,\varepsilon_{\mathcal{D}}=\mathrm{Ker}\,\mathcal{D}$. In particular, $\mathcal{D}\varepsilon_{\mathcal{D}}=0$.
Proof.
The module $\mathrm{Span}\,\varepsilon_{\mathcal{D}}$ consists of sequences of the form $a(\varepsilon_{\mathcal{D},i})$, where $a\in((\mathrm{Ker}\,{\mathcal{D}})^{*})^{*}=\mathrm{Ker}\,\mathcal{D}$. But $a(\varepsilon_{\mathcal{D},i})=\varepsilon_{\mathcal{D},i}(a)=a_{i}$, so $\mathrm{Span}\,\varepsilon_{\mathcal{D}}$ is precisely $\mathrm{Ker}\,\mathcal{D}$. In particular, by Proposition 5.5, we have $\mathcal{D}\varepsilon_{\mathcal{D}}=0$.∎
Corollary 5.9
For any $i\in\mathbb{Z}$, the list $\varepsilon_{\mathcal{D},i},\dots,\varepsilon_{\mathcal{D},i+d-1}$ is a basis of $(\mathrm{Ker}\,\,\mathcal{D})^{*}$.
Proof.
Apply Proposition 5.6. ∎
It follows that the sequence of points $P_{i}=R\varepsilon_{\mathcal{D},i}$ is a polygon in $\mathbb{P}((\mathrm{Ker}\,\,\mathcal{D})^{*})$. Thus, every properly bounded difference operator gives rise to a polygon in the projectivization of the dual to its kernel. This gives a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Denote the collection of properly bounded, degree $d$ left difference operators up to multiplication on the left and right by bi-infinite sequences of units as ${R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$.
Proposition 5.10
The map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ factors through a map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
Proof.
We need to check that if we take a properly bounded left difference operator $\mathcal{D}$, and multiply it on the left or right by a bi-infinite sequence of units in $R$, then the image of the resulting operator under $\bar{\varphi}$ is in the same projective equivalence class as the image of $\mathcal{D}$. Consider $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ and a bi-infinite sequence $c\in(\mathbb{R}^{\mathbb{Z}})^{*}$ of units in $R$. Note that multiplication on the left $\mathcal{D}\mapsto c\mathcal{D}$ does not change the kernel and hence the image under $\bar{\varphi}$. So, it suffices to establish projective equivalence of the polygons $\bar{\varphi}(\mathcal{D}c)$ and $\bar{\varphi}(\mathcal{D})$. Clearly, left multiplication by $c$ gives a right $R$-module isomorphism $\lambda_{c}\colon\mathrm{Ker}\,\mathcal{D}c\to\mathrm{Ker}\,\mathcal{D}$. Consider the associated left module isomorphism $\lambda_{c}^{*}\colon(\mathrm{Ker}\,\mathcal{D}c)^{*}\to(\mathrm{Ker}\,\mathcal{D})^{*}$. As before, let $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ be defined by $\varepsilon_{\mathcal{D}c,i}(a)=a_{i}$, and let $a\in\mathrm{Ker}\,\mathcal{D}$. Then
$\lambda_{c}^{*}(\varepsilon_{\mathcal{D}c,i})(a)=\varepsilon_{\mathcal{D}c,i}(\lambda_{c}(a))=\varepsilon_{\mathcal{D}c,i}(ca)=c_{i}a_{i}=c_{i}\varepsilon_{\mathcal{D},i}(a).$
So, $\lambda_{c}^{*}$ maps $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ to a scalar multiple of $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Therefore, the associated projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D}c)^{*})\to\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ maps the $R\varepsilon_{\mathcal{D}c,i}$ to $R\varepsilon_{\mathcal{D},i}$, as needed.∎
To establish injectivity, we start with the following.
Lemma 5.11
Suppose that $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ are properly bounded (left) difference operators of degree $d$. Assume also that $\mathrm{Ker}\,\mathcal{D}=\mathrm{Ker}\,\mathcal{D}^{\prime}$. Then $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$ for some $c^{\prime}\in(R^{\mathbb{Z}})^{*}$.
Proof.
Since $\mathcal{D}^{\prime}$ and $\mathcal{D}$ are properly bounded of the same degree $d$, there is a sequence of units $c^{\prime}\in(R^{\mathbb{Z}})^{*}$ such that $\mathcal{R}:=\mathcal{D}^{\prime}-c^{\prime}\mathcal{D}$ has degree at most $d-1$. Note that since $\mathrm{Ker}\,\mathcal{D}^{\prime}=\mathrm{Ker}\,\mathcal{D}$, we have $\mathrm{Ker}\,\mathcal{R}\supset\mathrm{Ker}\,\mathcal{D}$. So, by Proposition 5.4, it follows that $\mathcal{R}=0$, meaning that $\mathcal{D}^{\prime}=c^{\prime}\mathcal{D}$, as desired. ∎
Proposition 5.12
The map $\varphi:{R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is injective.
Proof.
Consider two properly bounded left difference operators $\mathcal{D},\mathcal{D}^{\prime}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\overline{\varphi}(\mathcal{D})=\overline{\varphi}(\mathcal{D}^{\prime})$. We need to show that $\mathcal{D}$ and $\mathcal{D}^{\prime}$ belong to the same class in the quotient ${R^{\mathbb{Z}}[T]_{d}^{pb}}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$.
The equality $\overline{\varphi}(\mathcal{D})=\overline{\varphi}(\mathcal{D}^{\prime})$ means that the polygons associated with $\mathcal{D}$, $\mathcal{D}^{\prime}$ are projectively equivalent. That is, there is a left module isomorphism $\psi\colon(\mathrm{Ker}\,\mathcal{D}^{\prime})^{*}\to(\mathrm{Ker}\,\mathcal{D})^{*}$ such that $\psi(\varepsilon_{\mathcal{D}^{\prime},i})=c_{i}\varepsilon_{\mathcal{D},i}$ for some sequence of units $c_{i}\in R$. Consider the dual $\psi^{*}\colon\mathrm{Ker}\,\mathcal{D}\to\mathrm{Ker}\,\mathcal{D}^{\prime}$. For any $a\in\mathrm{Ker}\,\mathcal{D}$, we have
$\psi^{*}(a)_{i}=\varepsilon_{\mathcal{D}^{\prime},i}(\psi^{*}(a))=\psi(\varepsilon_{\mathcal{D}^{\prime},i})(a)=c_{i}\varepsilon_{\mathcal{D},i}(a)=c_{i}a_{i}.$
So, $\psi^{*}$ is left multiplication by $c$, and $\mathrm{Ker}\,\mathcal{D}^{\prime}=c\mathrm{Ker}\,\mathcal{D}$. Therefore, $\mathrm{Ker}\,\mathcal{D}^{\prime}c=\mathrm{Ker}\,\mathcal{D}$, and, by the above lemma, we have $\mathcal{D}^{\prime}c=c^{\prime}\mathcal{D}$, as needed. ∎
Now, to complete the proof of Theorem 5.7, it suffices to show that the mapping $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ is surjective. In other words, given a polygon $(P_{i})$ in $\mathbb{P}(M)$. where $M$ is a rank $d$ free left $R$-module, we need to find a difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that the associated polygon in $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ is projectively equivalent to $(P_{i})$.
Proposition 5.13
Suppose $v_{i}\in M$ are such that the sequence $P_{i}=Rv_{i}$ is a polygon. Then there exists a properly bounded degree $d$ difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\mathcal{D}v=0$.
Proof.
Since $P_{i}$ is a polygon, the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$. In particular, there exist $\alpha_{i,j}\in R$ such that
$v_{i+d}=\alpha_{0,i}v_{i}+...+\alpha_{d-1,i}v_{i+d-1}.$
In other words, for the difference operator
$\mathcal{D}:=\alpha_{0}+\dots+\alpha_{d-1}T^{i+d-1}-T^{d},$
we have $\mathcal{D}v=0$. Further, for any $i\in\mathbb{Z}$ we have
$M=R\langle v_{i+1},\dots,v_{i+d}\rangle=R\langle v_{i+1},\dots,\dots,v_{i+d-1},\alpha_{0,i}v_{i}\rangle=P_{i+1}\oplus\dots\oplus P_{i+d-1}\oplus R\alpha_{0,i}v_{i}.$
Separately,
$M=P_{i+1}\oplus\dots\oplus P_{i+d-1}\oplus P_{i},$
so, together, we have
$P_{i}=Rv_{i}=R\alpha_{0,i}v_{i},$
and thus $\alpha_{0,i}$ is left-invertible and hence a unit for any $i\in\mathbb{Z}$. Therefore, $\mathcal{D}$ is properly bounded, as desired. ∎
The following says that the difference operator $\mathcal{D}$ from Proposition 5.13 provides a right inverse for the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Thus, this map is surjective, and the proof of Theorem 5.7 is complete.
Proposition 5.14
Assume that $v_{i}\in M$, where $M$ is a rank $d$ free left $R$-module, is such that the sequence $Rv_{i}$ is a polygon. Let $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ be a properly bounded degree $d$ difference operator such that $\mathcal{D}v=0$. Then there is a projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ mapping the polygon associated with $\mathcal{D}$ to the polygon $(Rv_{i})$.
Proof.
Since $\mathcal{D}v=0$, and the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$, by Proposition 5.6 we have $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. This means that the image of the mapping $\psi\colon M^{*}\to\mathbb{R}^{\mathbb{Z}}$ given by $\xi\mapsto(\xi(v_{i}))$ is precisely $\mathrm{Ker}\,\mathcal{D}$. Furthermore, since $R$ is stably finite, and $M^{*}$, $\mathrm{Ker}\,\mathcal{D}$ are free modules of the same rank, it follows that $\psi\colon M^{*}\to\mathrm{Ker}\,\mathcal{D}$ is a right $R$-module isomorphism. Consider the dual $\psi^{*}\colon(\mathrm{Ker}\,\mathcal{D})^{*}\to M$. It is a left $R$-module isomorphism. Further, for any $\xi\in M^{*}$, we have
$\xi(\psi^{*}(\varepsilon_{\mathcal{D},i}))=\varepsilon_{\mathcal{D},i}(\psi(\xi))=\xi(v_{i}),$
so $\psi^{*}(\varepsilon_{\mathcal{D},i})=v_{i}$, meaning that the associated projective map $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ takes the polygon associated with $\mathcal{D}$ to the polygon $Rv_{i}$, as needed. ∎
6 Integrability of pentagram maps over rings
In this section, we show that pentagram maps over arbitrary stably finite rings can be viewed as integrable systems. As a first step, in Section 6.1, we reformulate the pentagram map in the language of difference operators. This is used in Section 6.3 to give a Lax formulation of the pentagram map. It is given in terms of pseudo-difference operators, which are introduced in Section 6.2.
6.1 Pentagram maps in terms of difference operators
By Theorem 5.7, polygons in the (left) projective plane $\mathbb{P}^{2}(R)$, considered up to projective transformation, are in one-to-one correspondence with elements of the set $R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$ of properly bounded left difference operators of degree three, up to left-right action. Accordingly, one can describe the pentagram map as a dynamical system on difference operators. In this section we give an intrinsic description of that dynamical system. For technical reasons, we consider the inverse pentagram map, rather than the pentagram map itself. It turns out that that this inverse, phrased in terms of difference operators, is described by a simple linear equation. In the case when the ground ring is a field, this result was established in [8].
In order to have well-defined dynamics on difference operators, we only consider polygons whose image under the inverse pentagram map is a polygon (so that the image corresponds to a difference operator). Denote the inverse pentagram map by $\Psi$, and let $\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$ be the set of projective equivalence classes of polygons in $\mathbb{P}^{2}(R)$ whose image under the inverse pentagram map is a polygon. That way, we have a map
$\Psi\colon\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}\to\mathfrak{P}(\mathbb{P}^{2}(R)).$
For a degree three difference operator of the form
$\mathcal{D}=a+bT+cT^{2}+dT^{3},\quad a,b,c,d\in R^{\mathbb{Z}},$
let us define
$\mathcal{D}_{+}:=a+bT\hskip 28.45274pt\mathcal{D}_{-}:=cT^{2}+dT^{3}.$ (6.1)
Definition 6.1
Two properly bounded left difference operators of degree three, $\mathcal{D},\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, are said to be related, denoted $\mathcal{D}\sim\widehat{\mathcal{D}}$, if
$\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}.$ (6.2)
We note that if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}\sim\alpha\widehat{\mathcal{D}}$ for any $\alpha\in(R^{\mathbb{Z}})^{*}$. Therefore, we have a left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$.
Proposition 6.2
If $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})\subset\mathrm{Ker}\,\widehat{\mathcal{D}}$.
Proof.
By adding $\widehat{\mathcal{D}}_{+}\mathcal{D}_{+}$ to both sides, the definition (6.2) of the relation $\mathcal{D}\sim\widehat{\mathcal{D}}$ can be rewritten as
$\widehat{\mathcal{D}}_{+}\mathcal{D}=\widehat{\mathcal{D}}\mathcal{D}_{+}.$
In particular, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then, for any $\alpha\in\mathrm{Ker}\,\mathcal{D}$, we have $\widehat{\mathcal{D}}(\mathcal{D}_{+}\alpha)=\widehat{\mathcal{D}}_{+}(\mathcal{D}\alpha)=0.\qed$
Definition 6.3
We say that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is regular if there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, and the homomorphism of right $R$-modules given by
$\mathcal{D}_{+}\mid_{\mathrm{Ker}\,\mathcal{D}}\colon\mathrm{Ker}\,\mathcal{D}\to\mathrm{Ker}\,\widehat{\mathcal{D}}$
is surjective (and hence bijective).
Proposition 6.4
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is regular. Then
  1. 1.
    The operator $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ exists and is unique up to the left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$.
  2. 2.
    Any operator in the same orbit of the left-right $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$ action on $R^{\mathbb{Z}}[T]_{3}^{pb}$ as $\mathcal{D}$ is also regular. Moreover, for two regular operators in the same orbit, their related operators are also in the same orbit.
Proof.
1. Existence of $\widehat{\mathcal{D}}$ with the property $\mathcal{D}\sim\widehat{\mathcal{D}}$ is a part of the definition of a regular operator. That $\widehat{\mathcal{D}}$ also satisfies $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For any other $\widehat{\mathcal{D}}^{\prime}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$, we have $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\supset\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. Furthermore, we know that the mapping $\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}\to R^{3}$, given by taking three consecutive entries of a bi-infinite sequence, is an isomorphism. The same holds for $\widehat{\mathcal{D}}$. Therefore, the embedding $\mathrm{Ker}\,\widehat{\mathcal{D}}\to\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}$ must be an isomorphism as well. So, by Lemma 5.11, we have $\widehat{\mathcal{D}}^{\prime}=\alpha\widehat{\mathcal{D}}$ for some $\alpha\in(R^{\mathbb{Z}})^{*}$.
2. Consider $\widehat{\mathcal{D}}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ and $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. Take an operator $\mathcal{D}^{\prime}=\alpha\mathcal{D}\beta$ in the same orbit as $\mathcal{D}$. Let $\widehat{\mathcal{D}}^{\prime}=\widehat{\mathcal{D}}\alpha^{-1}$. Then $\mathcal{D}^{\prime}\sim\widehat{\mathcal{D}}^{\prime}$. Furthermore,
$\mathcal{D}_{+}^{\prime}(\mathrm{Ker}\,\mathcal{D}^{\prime})=\alpha\mathcal{D}_{+}\beta(\beta^{-1}\mathrm{Ker}\,\mathcal{D})=\alpha\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\alpha\mathrm{Ker}\,\widehat{\mathcal{D}}=\mathrm{Ker}\,\widehat{\mathcal{D}}^{\prime}.$
The result follows. ∎
As a result, for a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, all its related operators constitute a single orbit of the $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$ action. Moreover, this orbit remains the same if we replace $\mathcal{D}$ by a left-right-equivalent operator. Denote by $R^{\mathbb{Z}}[T]_{3}^{reg}$ the set of regular operators. Then we have a map
$\psi\colon R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$
which takes an orbit of a regular operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ to the orbit of its related operator $\widehat{\mathcal{D}}\sim\mathcal{D}$ (this orbit is independent on the choice of a related operator, by Proposition 6.4).
Recall also that, by Theorem 5.7, we have a bijection $\varphi\colon R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the correspondence between difference operators (up to left-right equivalence) and polygons (up to projective equivalence).
Theorem 6.5.
  1. 1.
    We have
    $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))=\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}.$ (6.3)
    In other words, regular difference operators precisely correspond to polygons which are mapped to polygons under the inverse pentagram map.
  2. 2.
    The following diagram commutes.
    ${R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})}$${R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})}$${\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}}$${\mathfrak{P}(\mathbb{P}^{2}(R)).}$$\scriptstyle{\psi}$$\scriptstyle{\varphi}$$\scriptstyle{\varphi}$$\scriptstyle{\Psi}$ (6.4)
    In other words, rephrased in terms of difference operators, the (inverse) pentagram map is described by equation (6.2).
Proof.
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ is regular. We will show that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$. This will establish the inclusion (6.3), as well as commutativity of the diagram (6.4). Let $v_{i}:=\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Then, by construction of the map $\bar{\varphi}\colon R^{\mathbb{Z}}[T]_{3}^{pb}\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the polygon $(Rv_{i})$ represents the class of polygons $\bar{\varphi}(\mathcal{D})$ associated to $\mathcal{D}$. Furthermore, by Proposition 5.8, we have that $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. Therefore,
$\mathrm{Span}(\mathcal{D}_{+}v)=\mathcal{D}_{+}(\mathrm{Span}\,v)=\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}},$
where $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ is any operator such that $\mathcal{D}\sim\widehat{\mathcal{D}}$. So, if we set $\hat{v}:=\mathcal{D}_{+}v$, by Proposition 5.6 we have that $\hat{v}_{i},\hat{v}_{i+1},\hat{v}_{i+2}$ is a basis in $(\mathrm{Ker}\,\widehat{\mathcal{D}})^{*}$ for all $i$. In other words, the points $R\hat{v}_{i}$ form a polygon. At the same time, if $\mathcal{D}=a+bT+cT^{2}+dT^{3}$, then
$\hat{v}_{i}=a_{i}v_{i}+b_{i}v_{i+1}=-c_{i}v_{i+2}-d_{i}v_{i+3}.$
So, $\hat{v}_{i}\in(Rv_{i}\oplus Rv_{i+1})\cap(Rv_{i+2}\oplus Rv_{i+3})$, meaning that the point $R\hat{v}_{i}$ lies on the lines $Rv_{i}\oplus Rv_{i+1}$, $Rv_{i+2}\oplus Rv_{i+3}$ (those are indeed lines by Lemma 4.6) and hence coincides with their intersection point (which is unique also by Lemma 4.6). In other words, the points $R\hat{v}_{i}$ are the vertices of the image of the polygon $(Rv_{i})$ under the inverse pentagram map, which proves that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$.
Now, we show that $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))\supset\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$. Take a polygon $(Rv_{i})$, $v_{i}\in M$ (where $M$ is a rank three free left $R$-module), such that its image under the inverse pentagram map is a polygon. We need to show that the corresponding difference operator $\mathcal{D}$ satisfying $\mathcal{D}v=0$ is regular. Arguing as above, we see that the image of the polygon $(Rv_{i})$ under the inverse pentagram map is given by the points $R\hat{v}_{i}$, where $\hat{v}:=\mathcal{D}_{+}v$. Since this image is a polygon, there exists $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\widehat{\mathcal{D}}\hat{v}=0$. Let’s show that $\mathcal{D}\sim\widehat{\mathcal{D}}$ and, furthermore, $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$. For the latter, note that
$\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathcal{D}_{+}(\mathrm{Span}\,v)=\mathrm{Span}(\mathcal{D}_{+}v)=\mathrm{Span}\,\hat{v}=\mathrm{Ker}\,\widehat{\mathcal{D}}.$
As for the former, observe that $\mathcal{D}_{+}(\mathrm{Ker}\,\mathcal{D})=\mathrm{Ker}\,\widehat{\mathcal{D}}$ implies that $\mathrm{Ker}\,\mathcal{D}\subset\mathrm{Ker}\,\tilde{\mathcal{D}}$, where
$\tilde{\mathcal{D}}:=\widehat{\mathcal{D}}_{-}\mathcal{D}_{+}-\widehat{\mathcal{D}}_{+}\mathcal{D}_{-}.$
Therefore, $\mathrm{Ker}\,\mathcal{D}$ is also contained in the kernel of a degree two operator $\mathrm{Ker}\,T^{-2}\tilde{\mathcal{D}}$, which, by Proposition 5.4, implies $T^{-2}\tilde{\mathcal{D}}=0$, and hence $\mathcal{D}\sim\widehat{\mathcal{D}}$. ∎
6.2 Pseudo-difference operators
A (left) pseudo-difference operator over a ring $R$ is, basically, a formal Laurent series in terms of the left shift operator $T$ with coefficients in $R^{\mathbb{Z}}$. Formally, consider the subring $R^{\mathbb{Z}}[T,T^{-1}]$ of $\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by $R^{\mathbb{Z}}$, $T$, and $T^{-1}$. The ring $R^{\mathbb{Z}}[T,T^{-1}]$ consists of operators of the form $\sum\alpha_{i}T^{i}$, where $\alpha_{i}\in R^{\mathbb{Z}}$, and $\alpha_{i}\neq 0$ for only finitely many values of $i$. Consider a valuation-type function $\nu\colon R^{\mathbb{Z}}[T,T^{-1}]\to\mathbb{Z}_{+}\cup+\infty$ given by the degree of the lowest order term, where, by definition, $\nu(0)=+\infty$. Then $||\cdot||:=\exp(-\nu(\cdot))$ is a norm on $R^{\mathbb{Z}}[T,T^{-1}]$, in the sense of [9, Definition 2.11].
Definition 6.6
The ring $R^{\mathbb{Z}}((T))$ of (left) pseudo-difference operators over a ring $R$ is the completion of $R^{\mathbb{Z}}[T,T^{-1}]$ with respect to the norm |||||$\cdot$||| | ⋅ | |.
Any pseudo-difference operator $\mathcal{Q}\in R^{\mathbb{Z}}((T))$ can be uniquely represented as a sum of a series $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$ where $\alpha_{i}\in R^{\mathbb{Z}}$, $m\in\mathbb{Z}$, and convergence is relative to the non-Archimedean norm.
Proposition 6.7
Suppose $\mathcal{Q}=\sum_{i=1}^{\infty}\alpha_{i}T^{i}$ is a power series in $T$ without a free term. Then $1+\mathcal{Q}$ is a unit in $R^{\mathbb{Z}}((T))$.
Proof.
We have $(1+\mathcal{Q})^{-1}=\sum_{i=0}^{\infty}(-\mathcal{Q})^{i}$. The series converges since the ring $R^{\mathbb{Z}}((T))$ is complete, and $||\mathcal{Q}||\leq e^{-1}$. ∎
Corollary 6.8
Let $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$. If $\alpha_{m}$ is a unit in $R^{\mathbb{Z}}$, then $\mathcal{Q}$ a unit in $R^{\mathbb{Z}}((T))$.
Proof.
We have $\mathcal{Q}=\alpha_{m}T^{m}(1+\mathcal{R})$ where $\mathcal{R}$ is a power series in $T$ without a free term. So, $\mathcal{Q}$ is a product of units and hence a unit. ∎
Remark 6.9
Note that $R^{\mathbb{Z}}((T))$ is not a subring of $\mathrm{End}_{R}(R^{\mathbb{Z}})$, as the action of a general pseudo-difference operator on a bi-infinite sequence is not well-defined. However, a pseudo-difference operator still defines an endomorphism of the right $R$-module of eventually vanishing sequences, i.e., bi-infinite sequences $\alpha_{i}\in R$ such that $\alpha_{i}=0$ for large enough $i$.
6.3 Lax representation for pentagram maps over rings
Given a projective equivalence class of polygons in $\mathbb{P}^{2}(R)$, consider the associated properly bounded difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, and let $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}\in R^{\mathbb{Z}}((T))$ (this is well-defined by Corollary 6.8). Note that since $\mathcal{D}$ is defined up to the left-right action of $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$, the operator $\mathcal{Q}$ is defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}((T))$. Hence, we have a map $\mathfrak{P}(\mathbb{P}^{2}(R))\to R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*},$ taking a projective equivalence class of polygons to the conjugacy class of the associated operator $\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$. We claim that this map provides a Lax representation of the pentagram map in the sense polygons related by the pentagram map are being mapped to conjugate elements of $R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*}$.
Proposition 6.10
Suppose $(P_{i})$ is a polygon in $\mathbb{P}^{2}(R)$ whose image $(\tilde{P}_{i})$ is also a polygon. Then the associated operators $\mathcal{Q}=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$ and $\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}$ (defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$) are conjugate in $R^{\mathbb{Z}}((T))$.
Proof.
By Theorem 6.5, the operators $\mathcal{D}$ and $\tilde{\mathcal{D}}$ can be chosen so that
$\mathcal{D}_{+}\tilde{\mathcal{D}}_{-}=\mathcal{D}_{-}\tilde{\mathcal{D}}_{+}.$
But then
$\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}=\tilde{\mathcal{D}}_{+}^{-1}\mathcal{Q}\tilde{\mathcal{D}}_{+},$
as desired. ∎
7 Invariants of pentagram maps on closed polygons
Here we use the Lax form found in the previous section to give invariants (i.e., conserved quantities, or first integrals) of the pentagram map on closed polygons (i.e., polygons with periodic vertices). In the case of a field, these invariants are known to form a maximal Poisson-commuting family [14]. We expect that similar results can be obtained for much bigger classes of rings.
7.1 Closed polygons and super-periodic difference operators
Definition 7.1
A difference/pseudo-difference operator $\mathcal{D}=\sum\alpha_{i}T^{i}$ is said to be periodic, with period $n$, if all its coefficients $\alpha_{i}$ are $n$-periodic sequences. Equivalently, $\mathcal{D}$ is $n$-periodic if $\mathcal{D}T^{n}=T^{n}\mathcal{D}$. An $n$-periodic difference operator is said to be super-periodic if its kernel consists entirely of $n$-periodic sequences.
This class of difference operators is naturally acted upon the left and right by the subgroup of periodic bi-infinite sequences of units.
Proposition 7.2
Let $\mathcal{D}$ be a super-periodic left properly bounded difference operator with period $n$. Consider also $\alpha,\beta\in(\mathbb{R}^{\mathbb{Z}})^{*}$. Then, $\alpha\mathcal{D}\beta$ is super-periodic with period $n$ if and only if $\alpha$ and $\beta$ are $n$-periodic.
Proof.
Suppose first that $\alpha$ and $\beta$ are $n$-periodic. Then
$\mathrm{Ker}\,\alpha\mathcal{D}\beta=\beta^{-1}\mathrm{Ker}\,\mathcal{D},$
so all elements of $\mathrm{Ker}\,\alpha\mathcal{D}\beta$ are $n$-periodic.
Conversely, suppose that that $\alpha\mathcal{D}\beta$ is super-periodic with period $n$. Fix $i\in\mathbb{Z}$, and choose $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$ such that $a_{i}=1$. Since $\alpha\mathcal{D}\beta$ is super-periodic with period $n$, we also have $a_{i+n}=1$. Further, since $a\in\mathrm{Ker}\,\alpha\mathcal{D}\beta$, we have $\beta a\in\mathrm{Ker}\,\,\mathcal{D}$. Since $\mathcal{D}$ is super-periodic with period $n$, this gives $\beta_{i}a_{i}=\beta_{i+n}a_{i+n}$, i.e, $\beta_{i+n}=\beta_{i}$. Since $i$ was arbitrary, this means that $\beta$ is $n$-periodic. But then $\alpha$ is a product of $n$-periodic pseudo-difference operators:
$\alpha=\alpha\mathcal{D}\beta(\mathcal{D}\beta)^{-1},$
and hence $n$-periodic. ∎
This gives the following periodic version of Theorem 5.7.
Corollary 7.3
Let $R$ be a stably finite ring. Then, there is a natural one-to-one correspondence between the following sets:
  1. 1.
    Projective equivalence classes of closed $n$-gons in the left projective space $\mathbb{P}^{d-1}(R)$.
  2. 2.
    Properly bounded super-periodic left difference operators of degree $d$ and period $n$ with coefficients in $R$, up to multiplication on the left and right by $n$-periodic sequences of units in $R$.
Proof.
If $\mathcal{D}$ is a super-periodic operator with period $n$, then $\varepsilon_{\mathcal{D}}$ is an $n$-periodic sequence in $(\mathrm{Ker}\,\mathcal{D})^{*}$. So, the mapping $\bar{\phi}\colon R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$ takes super-periodic operator with period $n$ to closed polygons with the same period. Conversely, given a closed $n$-gon $(P_{i})$, we can choose spanning vectors $v_{i}\in P_{i}$ so that $v_{i+n}=v_{i}$, and then the operator $\mathcal{D}$ such that $\mathcal{D}v=0$ can be taken to be $n$-periodic. So, there is a surjective mapping the set of properly bounded super-periodic left difference operators of degree $d$ and period $n$ to the set of projective equivalence classes of closed $n$-gons . Furthermore, by Proposition 7.2 preimages of every class of polygons under this mapping are precisely the orbits of the left-right action of $n$-periodic sequences of units.∎
One also has a periodic version of Theorem 6.5. In particular, the diagram (6.4) still commutes, assuming that we replace all objects with their periodic counterparts. As a result, to any projective equivalence class of closed $n$-gons we can assign an $n$-periodic pseudo-difference operator $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$, defined up to conjugation by $n$-periodic sequences of units. Upon application of the pentagram map, the operator $\mathcal{Q}$ changes to one that is conjugate to $\mathcal{Q}$ within the ring of $n$-periodic pseudo-difference operators.
7.2 The trace of a periodic pseudo-difference operator
For a ring $R$, consider the additive subgroup $[R,R]\subset R$ generated over $\mathbb{Z}$ by all commutators $[a,b]:=ab-ba$. The Abelian group $R/[R,R]$ is known as the cyclic space of $R$, and the quotient map $R\to R/[R,R]$ as the universal trace.
Definition 7.4
For an $n$-periodic left pseudo-difference operator $\mathcal{Q}=\sum\alpha_{i}T^{i}\in R^{\mathbb{Z}}((T))$, its trace $\mathrm{Tr}\mathcal{Q}$ is an element of the cyclic space $R/[R,R]$ defined by
$\mathrm{Tr}\,\mathcal{Q}:=\sum_{i=1}^{n}\alpha_{0,i},$
where the right-hand side is understood modulo $[R,R]$.
Thus, the trace is a homomorphism of Abelian groups $\mathrm{Tr}\colon R^{\mathbb{Z}}((T))\to R/[R,R]$. The following shows it is indeed a trace (i.e., satisfies the cyclic property).
Lemma 7.5
Let $\mathcal{Q}$ and $\widehat{\mathcal{Q}}$ be two $n$-periodic left pseudo-difference operators. Then
$\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}.$
Proof.
We need to show that $\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})=0$. Note that the left-hand side is bilinear over $\mathbb{Z}$. Furthermore, if we endow the cyclic space with discrete topology, then the trace function becomes continuous, so the operation $\mathcal{Q},\widehat{\mathcal{Q}}\mapsto\mathrm{Tr}(\mathcal{Q}\widehat{\mathcal{Q}}-\widehat{\mathcal{Q}}\mathcal{Q})$ distributes over infinite sums. Hence, it suffices to verify the statement in the case $\mathcal{Q}=\alpha T^{k}$, $\widehat{\mathcal{Q}}=\beta T^{k}$, where $\alpha,\beta\in R^{\mathbb{Z}}$ are $n$-periodic. Clearly, $\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=0$ if $l\neq-k$, so suppose $l=-k$. Then
$\mathrm{Tr}\,\mathcal{Q}\widehat{\mathcal{Q}}=\mathrm{Tr}\,\alpha T^{k}\beta T^{-k}=\sum_{i=1}^{n}\alpha_{i}\beta_{i+k}.$
while
$\mathrm{Tr}\,\widehat{\mathcal{Q}}\mathcal{Q}=\mathrm{Tr}\,\beta T^{-k}\alpha T^{k}=\sum_{i=1}^{n}\beta_{i}\alpha_{i-k}.$
Modulo $[R,R]$, these two expressions are equal. ∎
7.3 Invariants of the pentagram map
As discussed above, to any projective equivalence class of closed $n$-gons we can assign an $n$-periodic pseudo-difference operator $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$, defined up to conjugation by $n$-periodic bi-infinite sequences of units. Upon application of the pentagram map, the operator $\mathcal{Q}$ changes to a conjugate one. Now, consider the functions on the space $\mathfrak{P}(\mathbb{P}^{2}(R))$ of projective equivalence classes of polygons defined by
$f_{ij}:=\mathrm{Tr}(T^{in}\mathcal{\mathcal{Q}}^{j}),\hskip 28.45274pti\in\mathbb{Z},j\in\mathbb{Z}_{\geq 0},$ (7.1)
where $\mathcal{Q}$ is defined as described above.
Proposition 7.6
  1. 1.
    The functions $f_{ij}$ are independent of the choice of $\mathcal{Q}$ and, hence, are well-defined functions on $\mathfrak{P}(\mathbb{P}^{2}(R))$.
  2. 2.
    The functions $f_{ij}$ are invariant under the pentagram map.
Proof.
Since $\mathcal{Q}$ is defined up to conjugation, and the action of the pentagram map is also given by conjugation, it suffices to show that the functions $\mathcal{Q}\mapsto\mathrm{Tr}(T^{in}\mathcal{\mathcal{Q}}^{j})$ are conjugation-invariant (central). Indeed,
$\mathrm{Tr}(T^{in}(\widehat{\mathcal{Q}}\mathcal{Q}\widehat{\mathcal{Q}}^{-1})^{j})=\mathrm{Tr}(T^{in}\widehat{\mathcal{Q}}\mathcal{Q}^{j}\widehat{\mathcal{Q}}^{-1})=\mathrm{Tr}(\widehat{\mathcal{Q}}T^{in}\mathcal{Q}^{j}\widehat{\mathcal{Q}}^{-1})=\mathrm{Tr}(T^{in}\mathcal{Q}^{j}),$
where in the second equality we used periodicity of $\widehat{\mathcal{Q}}$, and in the last equality Lemma 7.5. ∎